OpenStudy (anonymous):

What is the surface area of this square pyramid? A. 144 cm2 B. 108 cm2 C. 189 cm2 D. 252 cm2

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

4 of those triangles plus 1 base

OpenStudy (anonymous):

ok

OpenStudy (danjs):

4 x (1/2) x 6 x 9 = 2 x 6 x 9 = 108 for the triangles

OpenStudy (danjs):

4 of them , times 1/2 base length times height length

OpenStudy (anonymous):

k is the answer 108

OpenStudy (danjs):

no, you have to include the square base too

OpenStudy (anonymous):

189

OpenStudy (danjs):

Base = length x length = 6 * 6 = 36

OpenStudy (danjs):

Total = 4 triangles + base = 108 + 36

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so is it 144 a

OpenStudy (danjs):

yep

OpenStudy (anonymous):

help me with a few more like 2